# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ 7 Experiments
Figure 1: AID vs ITD for synthetic elastic-net. $t$ corresponds to the number of steps to find an approximate solution to the fixed point and the dashed vertical line is the step where the support is identified. AID-FP converges faster than ITD; note that after support identification there is a wide gap between the methods, as anticipated by our theoretical bounds. AID-CG does not converge in plot on the right, probably due to sensitivity to numerical errors.
@@
Elastic Net  Let $(X,y)\in\mathbb{R}^{n\times p}\times\mathbb{R}^{n}$ be a training regression dataset. The elastic net solution $w(\lambda)$ is the minimizer of the objective function $\frac{1}{n}\lVert Xw-y\rVert^{2}+\frac{\lambda_{1}}{2}\lVert w\rVert^{2}_{2}+\lambda_{2}\lVert w\rVert_{1}$, where $\lambda=(\lambda_{1},\lambda_{2})$ are the regularization hyperparameters.
Data Poisoning  We consider a data poisoning scenario similar to the one in (Xiao et al., 2015), where an attacker would like to corrupt part of the training dataset by adding noise in order to decrease the accuracy of an Elastic-net regularized logistic regression model after training. In particular, let $c$ be the number of classes and $(\tilde{X},\tilde{y})\in\mathbb{R}^{n^{\prime}\times p}\times[c]^{n^{\prime}}$ be the examples to corrupt while $(X,y)\in\mathbb{R}^{n\times p}\times[c]^{n}$ are the clean ones. Let also $\Gamma\in\mathbb{R}^{n^{\prime}\times p}$ represent the noise and define the data poisoning elastic net solution as $w(\Gamma)=\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}f(\Gamma,w)+\frac{\lambda_{1}}{2}\lVert w\rVert_{2}+\lambda_{2}\lVert w\rVert_{1}$, where $f(\Gamma,w)=\ell(Xw,y)+\ell((\tilde{X}+\Gamma)w,\tilde{y})$ and $\ell$ is the cross-entropy loss. A strategy to find $\Gamma$ would be by approximating an element of the conservative Jacobian-vector product $D_{w}(\Gamma)^{\top}y$ where $y$ is the gradient of the cross-entropy loss on an hold out set. This setting is of particular interest, since $\Gamma$ is high dimensional and hence zero-order methods like grid or random search are less appropriate. For both settings and all considered methods, we find an approximate solution $w_{t}(\lambda)$ always by iterating the contraction map which describes the iterates of the deterministic iterative soft-thresholding algorithm (see e.g. (Combettes & Wajs, 2005)). Although this may be inefficient in the stochastic setup, it yields a fairer comparison, since both the stochastic and deterministic algorithms will have the same $w_{t}(\lambda)$ as input. Additional details are in the appendix.
@@
We compare our stochastic method NSID (Algorithm 1) against AID-FP and the algorithm SID in (Grazzi et al., 2023). In particular, for NSID $\hat{T}$ corresponds to one step of gradient descent on a minibatch of training points, while $G$ is soft-thresholding. We implement SID by setting in NSID $G(u,\lambda)=u$ and using $\hat{\Phi}(u,\lambda,\xi)=G(\hat{T}(u,\lambda,\xi),\lambda)$ in place of $\hat{T}$. Note that although the theoretical convergence guarantee for SID do not hold due to $\hat{\Phi}$ being biased, the performance of SID still effectively measures the impact of such bias in practice.
@@
Figure 2: Stochastic implicit differentiation for elastic net (left) and data poisoning (right). Mean (solid line) and the geometric standard deviation (shaded region) of the approximation error over 10 runs. SID does not converge on elastic net for a specific choice of $\lambda$ and diverges in data poisoning (hence we do not report it), while NSID converges faster (at the beginning) than the Deterministic AID-FP. Note that decreasing step-sizes provide a favorable choice.
@@ 8 Conclusions
We established convergence guarantees for nonsmooth implicit differentiation methods. Leveraging the foundation laid by (Bolte et al., 2022), we developed tools facilitating the translation of results from the smooth case. Specifically, we provided non-asymptotic linear convergence rates for AID-FP and ITD, focusing on deviations from their smooth analogs. Additionally, we introduced NSID, a principled stochastic algorithm. Numerical experiments underscored the distinctive behaviors of AID-FP and ITD, along with the good performance of NSID, which may be useful in large scale bilevel optimization problems in the future.
@@
This supplementary material is organized as follows. In App. A we recall the notion of definable mappings. App. B gives some auxiliary results and proof of lemmas in the main body. In App. C we present the proof of Theorem 4.1. App. D gives the proof of  Theorem 5.4. In App. E we address bilevel optimization. Finally, App. F contains more information on the numerical experiments.
@@ Appendix A Definable Mappings
The concept of definable sets and functions is part of the so called tame geometry. Here we give just a very brief account (additional details can be found in (Bolte & Pauwels, 2021)). An $o$-minimal structure on $(\mathbb{R},+,\cdot)$ (‘o’ stands for ’ordinal’) is a collection of sets $\mathcal{O}=(\mathcal{O}_{p})_{p\in\mathbb{N}}$, where each $\mathcal{O}_{p}$ is a Boolean algebra444 meaning a nonempty family of subset of $\mathbb{R}^{p}$ which is stable by complementations and finite unions and intersections. on $\mathbb{R}^{p}$ containing the algebraic sets (zeros of polynomial functions in $p$ variables), which enjoys also additional compatibility properties. Subsets of $\mathbb{R}^{p}$ which belongs to an $o$-minimal structure $\mathcal{O}$ are called definable in $\mathcal{O}$ and set-valued mappings $F\colon\mathbb{R}^{d}\rightrightarrows\mathbb{R}^{p}$ are said definable in $\mathcal{O}$ if their graphs (as a subset of $\mathbb{R}^{d+p}$) is definable in $\mathcal{O}$.
@@ Appendix B Auxiliary Lemmas
    Suppose that $p=p_{1}+p_{2}$ and set, for $k=1,2$
    $\mathcal{A}_{k}$ $=\mathrm{pr}_{k}(\mathcal{A})=\{A_{k}\in\mathbb{R}^{n\times p_{1}}\,|\,[A_{1},A_{2}]\in\mathcal{A}\}$
    $\mathcal{B}_{k}$ $=\mathrm{pr}_{k}(\mathcal{A})=\{B_{k}\in\mathbb{R}^{n\times p_{1}}\,|\,[B_{1},B_{2}]\in\mathcal{A}\},$
    where $\mathrm{pr}_{1}\colon\mathbb{R}^{d\times p}\to\mathbb{R}^{d\times p_{1}}$ and $\mathrm{pr}_{2}\colon\mathbb{R}^{d\times p}\to\mathbb{R}^{d\times p_{2}}$ are the canonical projections. Then $\operatorname{\it{e}}(\mathcal{A}_{k},\mathcal{B}_{k})\leq\operatorname{\it{e}}(\mathcal{A},\mathcal{B})$.
@@
which in turns holds for every $A_{1}\in\mathcal{A}_{1}$. Thus, taking the supremum in $A_{1}\in\mathcal{A}_{1}$ the statement follows with $k=1$. The other case is proved in the same manner. ∎
@@
Let $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a continuous selection of the definable Lipschitz smooth mappings $F_{1},\dots,F_{m}\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$. Let $L_{i}$ be the Lipschitz constant of $\partial F_{i}$ and set $L=\max_{1\leq i\leq m}L_{i}$. Then, for any $x\in U$ there exists $R_{x}>0$ such that
[…]
@@
$g\colon\left]0,+\infty\right[\rightrightarrows[m]\quad\text{such that}\ g(r)=I_{F}(B_{r}(x)),$
where $B_{r}(x)$ is the closed ball of radius $r>0$ centered at $x$. Now, we note that $g$ is the composition of the maps
$\varphi\colon\left]0,+\infty\right[\rightrightarrows\mathbb{R}^{p}\colon r\rightrightarrows B_{r}(x),\quad\text{and}\quad I_{F}\colon\mathbb{R}^{p}\rightrightarrows[m].$
The first one is clearly semialgebraic and hence definable and the second map is definable by definition (since the $F_{i}$’s are definable, it is easy to see that $F$ is definable if and only if $I_{F}$ is definable). Thus, being $g$ composition of definable set-valued mappings it is definable. Then, for every $I\subset[m]$, we have that the set $[g=I]=\{r\in\left]0,+\infty\right[\,|\,g(r)=I\}$ is definable and setting $\mathcal{J}=\{g(r)\,|\,r\in\left]0,+\infty\right[\}\subset 2^{[m]}$, we have that $([g=I])_{I\in\mathcal{J}}$, is a finite partition of $\left]0,+\infty\right[$ made of definable sets of the real line. Thus, each one of them must be finite unions of disjoints intervals, which shows that $g$ is piecewise constant. It follows that there exists $R_{x}>0$ and $I\subset[m]$ such that for every $r\in\left]0,R_{x}\right]$ $g(r)=I$. The proof continues as in Lemma 4 in (Bolte et al., 2022). ∎
@@
Let $x\in U$. Let $\Delta_{m}=\{\alpha\in\mathbb{R}_{+}^{m}\,|\,\sum_{i=1}^{m}\alpha_{i}=1\}$ be the unit simplex of $\mathbb{R}^{m}$ and $\Delta^{x}_{m}=\{\alpha\in\Delta_{m}\,|\,\forall\,i\in[m]\setminus I_{F}(x)\colon\alpha_{i}=0\}$ (which is essentially the unit simplex of $\mathbb{R}^{I_{F}(x)}$). Set $\mathcal{A}=\mathrm{co}(\{\partial F_{i}(x)\,|\,i\in I_{F}(x^{\prime})\})$. Then
@@
$(1)$ $=\sup_{\alpha\in\Delta_{m}^{x^{\prime}}}\inf_{\beta\in\Delta_{m}^{x^{\prime}}}\Big{\lVert}\sum_{i\in I(x^{\prime})}\alpha_{i}F^{\prime}_{i}(x^{\prime})-\sum_{i\in I(x^{\prime})}\beta_{i}F^{\prime}_{i}(x)\Big{\rVert}$
$\leq\sup_{\alpha\in\Delta_{m}^{x^{\prime}}}\Big{\lVert}\sum_{i\in I(x^{\prime})}\alpha_{i}(F^{\prime}_{i}(x^{\prime})-F^{\prime}_{i}(x))\Big{\rVert}$
$\leq\sup_{\alpha\in\Delta_{m}^{x^{\prime}}}\sum_{i\in I(x^{\prime})}\alpha_{i}\lVert F^{\prime}_{i}(x^{\prime})-F^{\prime}_{i}(x)\rVert\leq\sup_{\alpha\in\Delta_{m}^{x^{\prime}}}\sum_{i\in I(x^{\prime})}\alpha_{i}L\lVert x-x^{\prime}\rVert=L\lVert x-x^{\prime}\rVert.$
@@
$(2)$ $=\sup_{\alpha\in\Delta_{m}^{x^{\prime}}}\inf_{\beta\in\Delta_{m}^{x}}\Big{\lVert}\sum_{i\in I(x^{\prime})}\alpha_{i}F^{\prime}_{i}(x)-\sum_{i\in I(x)}\beta_{i}F^{\prime}_{i}(x)\Big{\rVert}=\sup_{\alpha\in\Delta_{m}^{x^{\prime}}}\inf_{\beta\in\Delta_{m}^{x}}\Big{\lVert}\sum_{i=1}^{m}(\alpha_{i}-\beta_{i})F^{\prime}_{i}(x)\Big{\rVert}$
$\leq\sup_{\alpha\in\Delta_{m}}\inf_{\beta\in\Delta_{m}^{x}}\Big{\lVert}\sum_{i=1}^{m}(\alpha_{i}-\beta_{i})F^{\prime}_{i}(x)\Big{\rVert}=:(*).$
@@
$\varphi(\alpha,\beta)=\Big{\lVert}\sum_{i=1}^{m}(\alpha_{i}-\beta_{i})F^{\prime}_{i}(x)\Big{\rVert}+\iota_{\Delta_{m}}(\alpha)+\iota_{\Delta_{m}^{x}}(\beta)$
is jointly convex, hence $\alpha\mapsto\inf_{\beta}\varphi(\alpha,\beta)$ is convex and its maximum is achieved at the vertices of $\Delta_{m}$. Thus, if we set $e_{i}=(\delta_{j}^{i})_{1\leq j\leq m}$ the canonical basis of $\mathbb{R}^{m}$, we have
(*(*)( * ) $=\max_{1\leq i\leq m}\inf_{\beta\in\Delta_{m}^{x}}\Big{\lVert}\sum_{j=1}^{m}(\delta^{i}_{j}-\beta_{j})F^{\prime}_{j}(x)\Big{\rVert}=\max_{1\leq i\leq m}\inf_{\beta\in\Delta_{m}^{x}}\Big{\lVert}F^{\prime}_{i}(x)-\sum_{j=1}^{m}\beta_{j}F^{\prime}_{j}(x)\Big{\rVert}$
$\leq\max_{1\leq i\leq m}\inf_{j\in I(x)}\Big{\lVert}F^{\prime}_{i}(x)-F^{\prime}_{j}(x)\Big{\rVert}=M_{x}.$
@@
Proof of Lemma 3.2.
It follows from Lemma 2.5 that the set-valued mapping
$(u,\lambda)\rightrightarrows D_{\Phi}^{s}(u,\lambda)=\operatorname{co}(\{\Phi^{\prime}_{i}(u,\lambda)\,|\,\in I_{\Phi}(u,\lambda)\})$
is a conservative derivative of $\Phi$ and for every $\lambda\in\Lambda$ there exists $R_{\lambda}>0$ (attached to $(w(\lambda),\lambda)$) such that
$\operatorname{\it{e}}(D_{\Phi}^{s}(u,\lambda),D_{\Phi}^{s}(w(\lambda),\lambda))\leq C_{\lambda}(u)\lVert u-w(\lambda)\rVert$ (22)
where
$C_{\lambda}(u):=\begin{cases}L&\text{if }\lVert u-w(\lambda)\rVert\leq R_{\lambda}\\ L+M_{\lambda}/R_{\lambda}&\text{otherwise}\end{cases}$
and $M_{\lambda}:=\max\limits_{i\in[m]}\min\limits_{j\in I_{\Phi}(w(\lambda),\lambda)}\lVert\Phi^{\prime}_{i}(w(\lambda),\lambda)-\Phi^{\prime}_{j}(w(\lambda),\lambda)\rVert$. Now, as already noted in the paper, it follows from Lemma 3 and Lemma 4 in (Bolte & Pauwels, 2021) that we can define a new conservative derivative of $\Phi$ as follows
$D_{\Phi}(u,\lambda)=D_{\Phi,1}^{s}(u,\lambda)\times D_{\Phi,2}^{s}(u,\lambda)$
and clearly $D_{\Phi,k}(u,\lambda)=D_{\Phi,k}^{s}(u,\lambda)=\operatorname{co}(\{\partial_{k}\Phi_{i}(u,\lambda)\,|\,i\in I_{\Phi}(u,\lambda)\})$ for $k=1,2$. Therefore the statement follows from (22) and Lemma B.1(vi). ∎
@@
$\lVert(I-D_{\Phi,1}(u,\lambda))^{-1}\rVert\leq\frac{1}{1-q},\qquad\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup}\leq\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}\leq\frac{\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}}{1-q}.$

### v2

@@ 7 Experiments
Refer to captionRefer to caption
Figure 1: AID vs ITD for synthetic elastic-net. $t$ corresponds to the number of steps to find an approximate fixed point and the dashed vertical line is the step where the support is identified. AID-FP converges faster than ITD; note that after support identification there is a wide gap between the methods, as anticipated by our theoretical bounds. AID-CG does not converge in plot on the right, probably due to sensitivity to numerical errors.
@@
Elastic Net  Let $(X,y)\in\mathbb{R}^{n\times d}\times\mathbb{R}^{n}$ be a training regression dataset. The elastic net solution $w(\lambda)$ is the minimizer of the objective function $\frac{1}{n}\lVert Xw-y\rVert^{2}+\frac{\lambda_{1}}{2}\lVert w\rVert^{2}_{2}+\lambda_{2}\lVert w\rVert_{1}$, where $\lambda=(\lambda_{1},\lambda_{2})$ are the regularization hyperparameters.
Data Poisoning  We consider a data poisoning scenario similar to the one in (Xiao et al., 2015), where an attacker would like to corrupt part of the training dataset by adding noise in order to decrease the accuracy of an Elastic-net regularized logistic regression model after training. In particular, let $c$ be the number of classes and $(\tilde{X},\tilde{y})\in\mathbb{R}^{n^{\prime}\times d}\times[c]^{n^{\prime}}$ be the examples to corrupt while $(X,y)\in\mathbb{R}^{n\times d}\times[c]^{n}$ are the clean ones. Let also $\Gamma\in\mathbb{R}^{n^{\prime}\times d}$ represent the noise and define the data poisoning elastic net solution as $w(\Gamma)=\operatorname*{arg\,min}_{w\in\mathbb{R}^{d}}f(\Gamma,w)+\frac{\lambda_{1}}{2}\lVert w\rVert_{2}+\lambda_{2}\lVert w\rVert_{1}$, where $f(\Gamma,w)=\ell(Xw,y)+\ell((\tilde{X}+\Gamma)w,\tilde{y})$ and $\ell$ is the cross-entropy loss. A strategy to find $\Gamma$ would be by approximating an element of the conservative Jacobian-vector product $D_{w}(\Gamma)^{\top}y$ where $y$ is the gradient of the cross-entropy loss on an hold out set. This setting is of particular interest, since $\Gamma$ is high dimensional and hence zero-order methods like grid or random search are less appropriate. For both settings and all considered methods, we find an approximate solution $w_{t}(\lambda)$ always by iterating the contraction map which describes the iterates of the deterministic iterative soft-thresholding algorithm (see e.g., (Combettes & Wajs, 2005)). Although this may be inefficient in the stochastic setup, it yields a fairer comparison, since both the stochastic and deterministic algorithms will have the same $w_{t}(\lambda)$ as input. Additional details are in the appendix.
@@
We compare our stochastic method NSID (Algorithm 1) against AID-FP and the algorithm SID in (Grazzi et al., 2023). In particular, for NSID $\hat{T}_{x}$ corresponds to one step of gradient descent on a minibatch of training points, while $G$ is soft-thresholding. We implement SID by setting in NSID $G(u,\lambda)=u$ and using $\hat{\Phi}_{\xi}(u,\lambda)=G(\hat{T}_{\xi}(u,\lambda),\lambda)$ in place of $\hat{T}_{\xi}$. Note that although the theoretical convergence guarantee for SID do not hold due to $\hat{\Phi}_{\xi}$ being biased, the performance of SID still effectively measures the impact of such bias in practice.
@@
Refer to captionRefer to caption
Figure 2: Stochastic implicit differentiation for elastic net (left) and data poisoning (right) with constant (const) and decreasing (dec) step sizes. Mean (solid line) and the geometric standard deviation (shaded region) of the approximation error over 10 runs. SID does not converge on elastic net for a specific choice of $\lambda$ and diverges in data poisoning (hence we do not report it), while NSID converges faster (at the beginning) than the Deterministic AID-FP. Note that decreasing step-sizes provide a favorable choice.
@@ 8 Conclusions
We established convergence guarantees for nonsmooth implicit differentiation methods. Leveraging the foundation laid by (Bolte et al., 2022), we developed tools facilitating the translation of results from the smooth case. This allowed us to provided non-asymptotic linear convergence rates for AID-FP and ITD, focusing on deviations from their smooth analogs. Additionally, we introduced NSID, a principled stochastic algorithm. Numerical experiments underscored the distinctive behaviors of AID-FP and ITD, along with the good performance of NSID, which may be useful in large scale bilevel optimization problems in the future. Despite our results, establishing rates for solving nonsmooth bilevel problems is still challenging and we leave it for future work.
@@
This supplementary material is organized as follows. In App. A we recall the notion of definable mappings. App. B gives some auxiliary results and proof of lemmas in the main body. In App. C we present the proof of Theorem 4.1. App. D gives the proof of  Theorems 5.5 and 5.6. In App. E we address bilevel optimization. Finally, App. F contains more information on the numerical experiments.
@@ Appendix A Definable Mappings
The concept of definable sets and functions is part of the so called tame geometry. Here we give just a very brief account (additional details can be found in (Bolte & Pauwels, 2021)). An $o$-minimal structure on $(\mathbb{R},+,\cdot)$ (‘o’ stands for ’ordinal’) is a collection of sets $\mathcal{O}=(\mathcal{O}_{p})_{p\in\mathbb{N}}$ such that, for each $p\in\mathbb{N}$,
  1. (i)
    $\mathcal{O}_{p}$ is a Boolean algebra, meaning a nonempty family of subset of $\mathbb{R}^{p}$ which is stable by complementations and finite unions and intersections. Moreover, it contains the algebraic sets, that is, the sets of zeros of polynomial functions in $p$ variables.
  2. (ii)
    $\mathcal{O}_{1}$ is made exactly of finite unions of intervals.
  3. (iii)
    $A\in\mathcal{O}_{p}\Rightarrow\ A\times\mathbb{R},\mathbb{R}\times A\in\mathcal{O}_{p+1}$
  4. (iv)
    if $\pi_{p}\colon\mathbb{R}^{p+1}\to\mathbb{R}^{p}$ is the canonical projection onto the first $p$ components, then $A\in\mathcal{O}_{p+1}\ \Rightarrow\ \pi_{p}(A)\in\mathcal{O}_{p}$;
Subsets of $\mathbb{R}^{p}$ which belongs to an $o$-minimal structure $\mathcal{O}$ are called definable in $\mathcal{O}$ and set-valued mappings $F\colon\mathbb{R}^{d}\rightrightarrows\mathbb{R}^{p}$ are said definable in $\mathcal{O}$ if their graphs (as a subset of $\mathbb{R}^{d+p}$) is definable in $\mathcal{O}$.
@@ Appendix B Auxiliary Lemmas
    Suppose that $p=p_{1}+p_{2}$ and set, for $k=1,2$ $\mathrm{pr}_{1}\colon\mathbb{R}^{n\times p}\to\mathbb{R}^{n\times p_{k}}$ be the canonical projections and
    $\mathcal{A}_{k}$ $=\mathrm{pr}_{k}(\mathcal{A})=\{A_{k}\in\mathbb{R}^{n\times p_{k}}\,|\,[A_{1},A_{2}]\in\mathcal{A}\},\quad\mathcal{B}_{k}=\mathrm{pr}_{k}(\mathcal{B})=\{B_{k}\in\mathbb{R}^{n\times p_{k}}\,|\,[B_{1},B_{2}]\in\mathcal{A}\}.$
    Then $\operatorname{\it{e}}(\mathcal{A}_{k},\mathcal{B}_{k})\leq\operatorname{\it{e}}(\mathcal{A},\mathcal{B})$.
  7. (vii)
    Suppose that $p=p_{1}+p_{2}$. Then, for all $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{p_{1}\times p_{2}}$, we have
    $\lVert\mathcal{A}(\mathcal{X})\rVert_{\sup}\leq\lVert\mathcal{A}_{1}\rVert_{\sup}\lVert\mathcal{X}\rVert_{\sup}+\lVert\mathcal{A}_{2}\rVert_{\sup},$
    $\operatorname{\it{e}}(\mathcal{A}(\mathcal{X}),\mathcal{A}(\mathcal{\mathcal{Y}}))\leq\lVert\mathcal{A}_{1}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{X},\mathcal{Y}),\qquad\operatorname{\it{e}}(\mathcal{A}(\mathcal{X}),\mathcal{B}(\mathcal{X}))\leq(1+\lVert\mathcal{X}\rVert_{\sup})\operatorname{\it{e}}(\mathcal{A},\mathcal{B})$
    where we recall that $\mathcal{A}(\mathcal{X})=\{A_{1}X+A_{2}\,|\,[A_{1},A_{2}]\in\mathcal{A},X\in\mathcal{X}\}$.
@@
which in turns holds for every $A_{1}\in\mathcal{A}_{1}$. Thus, taking the supremum in $A_{1}\in\mathcal{A}_{1}$ the statement follows with $k=1$. The other case is proved in the same manner.
(vii): For the first inequality we have
$\lVert\mathcal{A}(\mathcal{X})\rVert_{\sup}$ $=\sup_{A\in\mathcal{A},X\in\mathcal{X}}\lVert A_{1}X+A_{2}\rVert$
$\leq\sup_{A\in\mathcal{A},X\in\mathcal{X}}\left(\lVert A_{1}\rVert\lVert X\rVert+\lVert A_{2}\rVert\right)$
$\leq\sup_{A\in\mathcal{A}}\lVert A_{1}\rVert\sup_{A\in\mathcal{X}}\lVert X\rVert+\sup_{A^{\prime}\in\mathcal{A}}\lVert A^{\prime}_{2}\rVert=\lVert\mathcal{A}_{1}\rVert_{\sup}\lVert\mathcal{X}\rVert_{\sup}+\lVert\mathcal{A}_{2}\rVert_{\sup}.$
For the second inequality we have
$\operatorname{\it{e}}(\mathcal{A}(\mathcal{X}),\mathcal{A}(\mathcal{Y}))$ $=\sup_{A\in\mathcal{A},X\in\mathcal{X}}\inf_{A^{\prime}\in\mathcal{A},Y\in\mathcal{Y}}\lVert A_{1}X-A_{2}-A^{\prime}_{1}Y+A^{\prime}_{2}\rVert$
$\leq\sup_{A\in\mathcal{A},X\in\mathcal{X}}\inf_{Y\in\mathcal{Y}}\lVert A_{1}(X-Y)\rVert$
$\leq\sup_{A\in\mathcal{A}}\lVert A_{1}\rVert\sup_{X\in\mathcal{X}}\inf_{Y\in\mathcal{Y}}\lVert X-Y\rVert=\lVert\mathcal{A}_{1}\rVert_{\sup}\operatorname{\it{e}}(\mathcal{X},\mathcal{Y}).$
For the third inequality we have
$\operatorname{\it{e}}(\mathcal{A}(\mathcal{X}),\mathcal{B}(\mathcal{X}))$ $=\sup_{A\in\mathcal{A},X\in\mathcal{X}}\inf_{B\in\mathcal{B},X^{\prime}\in\mathcal{X}}\lVert A_{1}X-A_{2}-B_{1}X^{\prime}+B_{2}\rVert$
$\leq\sup_{A\in\mathcal{A},X\in\mathcal{X}}\inf_{B\in\mathcal{B}}\lVert(A_{1}-B_{1})X-A_{2}+B_{2}\rVert$
$\leq\sup_{A\in\mathcal{A},X\in\mathcal{X}}\inf_{B\in\mathcal{B}}\left(\lVert A_{1}-B_{1}\rVert\lVert X\rVert+\lVert A_{2}-B_{2}\rVert\right)$
$\leq\sup_{A\in\mathcal{A},X\in\mathcal{X}}\inf_{B\in\mathcal{B}}\left(\lVert A-B\rVert\lVert X\rVert+\lVert A-B\rVert\right)$
$\leq\sup_{X\in\mathcal{X}}(1+\lVert X\rVert)\sup_{A\in\mathcal{A}}\inf_{B\in\mathcal{B}}\lVert A-B\rVert=(1+\lVert\mathcal{X}\rVert_{\sup})\operatorname{\it{e}}(\mathcal{A},\mathcal{B}).$
The proof is complete. ∎
@@
Let $F\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$ be a continuous selection of the definable Lipschitz smooth mappings $F_{1},\dots,F_{r}\colon U\subset\mathbb{R}^{p}\to\mathbb{R}^{d}$. Let $L_{i}$ be the Lipschitz constant of $F^{\prime}_{i}$ and set $L=\max_{1\leq i\leq r}L_{i}$. Then, for any $x\in U$ there exists $R_{x}>0$ such that
$\forall\,x^{\prime}\in U\ \text{with}\ \lVert x^{\prime}-x\rVert\leq R_{x}\colon\ \operatorname{\it{e}}(D_{F}^{s}(x^{\prime}),D_{F}^{s}(x))\leq L\lVert x^{\prime}-x\rVert.$
@@
$g\colon\left]0,+\infty\right[\rightrightarrows[r]\quad\text{such that}\ g(\rho)=I_{F}(B_{\rho}(x)),$
where $B_{\rho}(x)$ is the closed ball of radius $\rho>0$ centered at $x$. Now, we note that $g$ is the composition of the maps
$\varphi\colon\left]0,+\infty\right[\rightrightarrows\mathbb{R}^{p}\colon\rho\rightrightarrows B_{\rho}(x),\quad\text{and}\quad I_{F}\colon\mathbb{R}^{p}\rightrightarrows[r].$
The first one is clearly semialgebraic and hence definable and the second map is definable by definition (since the $F_{i}$’s are definable, it is easy to see that $F$ is definable if and only if $I_{F}$ is definable). Thus, being $g$ composition of definable set-valued mappings it is definable. Then, for every $I\subset[r]$, we have that the set $[g=I]=\{\rho\in\left]0,+\infty\right[\,|\,g(\rho)=I\}$ is definable and setting $\mathcal{J}=\{g(\rho)\,|\,\rho\in\left]0,+\infty\right[\}\subset 2^{[r]}$, we have that $([g=I])_{I\in\mathcal{J}}$, is a finite partition of $\left]0,+\infty\right[$ made of definable sets of the real line. Thus, each one of them must be finite unions of disjoints intervals, which shows that $g$ is piecewise constant. It follows that there exists $R_{x}>0$ and $I\subset[r]$ such that for every $\rho\in\left]0,R_{x}\right]$ $g(\rho)=I$. The proof continues as in Lemma 4 in (Bolte et al., 2022). ∎
@@
Let $x\in U$. Let $\Delta_{r}=\{\alpha\in\mathbb{R}_{+}^{r}\,|\,\sum_{i=1}^{r}\alpha_{i}=1\}$ be the unit simplex of $\mathbb{R}^{r}$ and $\Delta^{x}_{r}=\{\alpha\in\Delta_{r}\,|\,\forall\,i\in[r]\setminus I_{F}(x)\colon\alpha_{i}=0\}$ (which is essentially the unit simplex of $\mathbb{R}^{I_{F}(x)}$). Set $\mathcal{A}=\mathrm{co}(\{\partial F_{i}(x)\,|\,i\in I_{F}(x^{\prime})\})$. Then, using the property of the excess in Lemma B.1(i)
@@
$(1)$ $=\sup_{\alpha\in\Delta_{r}^{x^{\prime}}}\inf_{\beta\in\Delta_{r}^{x^{\prime}}}\Big{\lVert}\sum_{i\in I(x^{\prime})}\alpha_{i}F^{\prime}_{i}(x^{\prime})-\sum_{i\in I(x^{\prime})}\beta_{i}F^{\prime}_{i}(x)\Big{\rVert}$
$\leq\sup_{\alpha\in\Delta_{r}^{x^{\prime}}}\Big{\lVert}\sum_{i\in I(x^{\prime})}\alpha_{i}(F^{\prime}_{i}(x^{\prime})-F^{\prime}_{i}(x))\Big{\rVert}$
$\leq\sup_{\alpha\in\Delta_{r}^{x^{\prime}}}\sum_{i\in I(x^{\prime})}\alpha_{i}\lVert F^{\prime}_{i}(x^{\prime})-F^{\prime}_{i}(x)\rVert\leq\sup_{\alpha\in\Delta_{r}^{x^{\prime}}}\sum_{i\in I(x^{\prime})}\alpha_{i}L\lVert x-x^{\prime}\rVert=L\lVert x-x^{\prime}\rVert.$
@@
$(2)$ $=\sup_{\alpha\in\Delta_{r}^{x^{\prime}}}\inf_{\beta\in\Delta_{r}^{x}}\Big{\lVert}\sum_{i\in I(x^{\prime})}\alpha_{i}F^{\prime}_{i}(x)-\sum_{i\in I(x)}\beta_{i}F^{\prime}_{i}(x)\Big{\rVert}=\sup_{\alpha\in\Delta_{r}^{x^{\prime}}}\inf_{\beta\in\Delta_{r}^{x}}\Big{\lVert}\sum_{i=1}^{r}(\alpha_{i}-\beta_{i})F^{\prime}_{i}(x)\Big{\rVert}$
$\leq\sup_{\alpha\in\Delta_{r}}\inf_{\beta\in\Delta_{r}^{x}}\Big{\lVert}\sum_{i=1}^{r}(\alpha_{i}-\beta_{i})F^{\prime}_{i}(x)\Big{\rVert}=:(*).$
@@
$\varphi(\alpha,\beta)=\Big{\lVert}\sum_{i=1}^{r}(\alpha_{i}-\beta_{i})F^{\prime}_{i}(x)\Big{\rVert}+\iota_{\Delta_{r}}(\alpha)+\iota_{\Delta_{r}^{x}}(\beta)$
is jointly convex, hence $\alpha\mapsto\inf_{\beta}\varphi(\alpha,\beta)$ is convex and its maximum is achieved at the vertices of $\Delta_{r}$. Thus, if we set $e_{i}=(\delta_{j}^{i})_{1\leq j\leq r}$ the canonical basis of $\mathbb{R}^{r}$, we have
((*)( ∗ ) $=\max_{1\leq i\leq r}\inf_{\beta\in\Delta_{r}^{x}}\Big{\lVert}\sum_{j=1}^{r}(\delta^{i}_{j}-\beta_{j})F^{\prime}_{j}(x)\Big{\rVert}=\max_{1\leq i\leq r}\inf_{\beta\in\Delta_{r}^{x}}\Big{\lVert}F^{\prime}_{i}(x)-\sum_{j=1}^{r}\beta_{j}F^{\prime}_{j}(x)\Big{\rVert}$
$\leq\max_{1\leq i\leq r}\inf_{j\in I(x)}\Big{\lVert}F^{\prime}_{i}(x)-F^{\prime}_{j}(x)\Big{\rVert}=M_{x}.$
@@
$\lVert(I-D_{\Phi,1}(u,\lambda))^{-1}\rVert_{\sup}\leq\frac{1}{1-q},\qquad\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup}\leq\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}\leq\frac{\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}}{1-q}.$